1 0
Add a comment Improve this question Transcribed image text
Answer #1

A)

The force R on the man will be in upward direction .

a) When a = 1.5 upwards

ma = R- mg

R = m(g+a)= 70 (9.8 + 1.5) =\mathbf{ 791 N}

b) When a = 1.5 downwards

ma = mg - R

R = m(g-a)= 70 (9.8 - 1.5) =\mathbf{ 581 N}

c) When velcoity is constant , a = 0

ma = R- mg = 0

R = m(g)= 70 (9.8) =\mathbf{ 686 N}

B) Similarly , The cables support the system consisting of man and elevator.

So, R = 4T

and M = 70 +1500 = 1570

a) When a = 1.5 upwards

Ma = 4T- Mg

T =\frac{ M(g+a)}{4}= \frac{1570 (9.8 + 1.5)}{4} =\mathbf{ 4435.25 N}

b) When a = 1.5 downwards

Ma = Mg - 4T

T=\frac{ M(g-a)}{4}= \frac{1570 (9.8 - 1.5)}{4} =\mathbf{ 3257.75 N}

c) When velcoity is constant , a = 0

Ma = 4T- Mg = 0

T =\frac{ Mg}{4}= \frac{1570 (9.8)}{4} =\mathbf{ 3846.5 N}
Answer #2

Hello Friend,

.com

Answer for first question

A.

a. N-m*g=m*a

and we get N = m (g+a) = 791 N

b. N-m*g =-m*a

N = m(g-a)=581 N

c.as velocity is constant , a= 0

that is N = m*g = 686 N

B.

4T